OpenStudy (anonymous):

At the park there is a pool shaped like a circle with diameter 20 yards . A ring-shaped path goes around the pool. Its width is 5 yards. We are going to give a new layer of coating to the path. If one gallon of coating can cover 5 yards, how many gallons of coating do we need? Note that coating comes only by the gallon, so the number of gallons must be a whole number. (Use the value 3.14 for pi.)

jimthompson5910 (jim_thompson5910):

We have 2 concentric circles 

jimthompson5910 (jim_thompson5910):

To find the area of the outer ring, follow these steps 1) Find the area of the larger circle (call it A) 2) Find the area of the smaller circle (call it B) 3) Subtract the two areas: A - B

OpenStudy (anonymous):

I found the area of the bigger circle which is 314 but how would i find the area of the smaller circle?

OpenStudy (anonymous):

Would I do (2.5^2)(3.14)?

jimthompson5910 (jim_thompson5910):

Keep in mind that the smaller circle has a diameter of 20 (so it has a radius of 10)

jimthompson5910 (jim_thompson5910):

Area of smaller circle: A = pi*r^2 A = pi*10^2 A = pi*100 A = 100pi A = 100(3.14) A = 314 So the area of the smaller circle is 314 square yards

OpenStudy (anonymous):

And the area of the larger circle would be found with what equation? (3.14)(2.5^2)?

jimthompson5910 (jim_thompson5910):

(3.14)(15)^2

jimthompson5910 (jim_thompson5910):

since the radius of the new larger circle is 15 yards

OpenStudy (anonymous):

So the ending result would be 392.5?

jimthompson5910 (jim_thompson5910):

good, now divide that by 5 (since 1 gallon of paint can do 5 square yards)

OpenStudy (anonymous):

78.5?

OpenStudy (anonymous):

so 79?

jimthompson5910 (jim_thompson5910):

good, the answer is 79

OpenStudy (anonymous):

Thank you so much!! :) :) That helped me a lot!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Do you know how to factor?

jimthompson5910 (jim_thompson5910):

I do, what's your question?